# Xpikeformer: Hybrid Analog-Digital Hardware Acceleration for Spiking Transformers

### v1

@@ VII-A3 Results
Fig. 11 presents the breakdown of Xpikeformer computational energy consumption. The majority of the computing energy is consumed by the AIMC engine, accounting for 73.5% of the total computational energy. The SSA engine contributes 18.7%, while other components, such as residual units, collectively account for 7.8%. A more detailed breakdown of the AIMC engine’s energy consumption is shown on the right side of Fig. 11. Notably, the ADC consumes about 2.0% of the AIMC energy. This efficiency is achieved by sharing ADCs across multiple read-out channels through multiplexing (Xpikeformer uses a sharing ratio of 8, balancing energy efficiency and latency), which significantly reduces the static power consumption that would otherwise be incurred by using dedicated ADCs for each channel. We also observe that the periphery circuits—decoders, multiplexers, switch matrices, buffers, etc., consume 86.0% of the AIMC energy. This is because ADC multiplexing requires a dedicated decoder and control logic, making these components relatively more energy-intensive than the crossbar itself. Additionally, the SA mapping strategy necessitates dedicated mapping control between the shared SRAM and local SA buffers. This overhead is unavoidable in this work due to the crossbar size limitations discussed in Section IV-A. Finally, the accumulation processes, including differential adders and LIF units, which are digital operations, contribute 12.0% of the AIMC energy.
In the two tasks tested, Xpikeformer achieves a 17.8–19.2$\times$ energy reduction compared to the SOTA digital ANN transformer accelerator (ANN-Digital), a 5.0–7.1$\times$ energy reduction compared to the hybrid implementation of ANN (ANN-Hybrid), and a 5.9–6.8$\times$ energy reduction compared to the fully digital implementation of spiking transformers on general-purpose devices (SNN-Digital (no SSA)).
VII-B Latency Evaluation
We compare the latency of Xpikeformer with two same-sized baseline implementations: an ANN transformer and a spiking transformer deployed on a GPU (Nvidia RTX A2000). We evaluate the runtime per image in Task 1. We adopt a spike encoding length $T=10$ for both Xpikeformer and the spiking transformer baseline.
Our observations show that Xpikeformer achieves a 1.53×1.53$\times$1.53 × speedup compared to the ANN GPU implementation, primarily due to the fast MVM conducted in-memory and the highly parallel and streamlined design of the SSA. Additionally, the spiking transformer exhibits poor time efficiency when deployed on GPU devices due to intrinsic precision mismatches and the lack of dedicated optimization. In contrast, Xpikeformer demonstrates a 12.0×12.0$\times$12.0 × speedup over the spiking transformer.
Figure 12: Speedup of Xpikeformer compared with (a) an ANN Transformer and (b) a spiking transformer deployed on GPU devices.
VII-C FPGA Implementation of SSA
We also implemented the proposed SSA tile on a lightweight FPGA (within Xilinx Zynq-7000 SoC). The latency and power consumption of the FPGA implementation were compared against CPU (Intel i7-12850HX) and GPU (Nvidia RTX A2000), as shown in Table IV. We obtain 48×48$\times$48 × better latency expending 15×15$\times$15 × lesser power than that of the GPU implementation. We also note that the SSA block on FPGA has 18×18$\times$18 × less latency while consuming 17×17$\times$17 × less power than the ANN GPU implementation.
TABLE IV: Comparison of hardware efficiency for a single attention head (CPU and GPU implementations) for different architectures. For SSA block, $T=10$.
Architecture – Device $f_{\text{clk}}$ (MHz) Latency (ms) Power (W)
ANN attention – CPU $2100$ $0.15$ $107.01$
ANN attention – GPU $562$ $0.06$ $26.13$
SSA – CPU $2100$ $2.672$ $65.54$
SSA – GPU $562$ $0.159$ $22.41$
SSA – FPGA $200$ $\mathbf{3.3\times 10^{-3}}$ $\mathbf{1.47}$
@@ VIII Conclusion
In this paper, we introduced Xpikeformer, a novel hybrid analog-digital hardware architecture designed to accelerate spiking neural network (SNN)-based transformer models. The architecture integrates AIMC for feedforward and fully connected layers and employs a stochastic spiking attention (SSA) engine to efficiently handle the attention mechanism. Our design mitigates the data communication overhead and computational inefficiencies typically associated with deploying spiking transformers on general-purpose platforms.
Xpikeformer demonstrates software-comparable inference accuracy, achieving 84.74% on the CIFAR-10 image classification task and maintaining accuracy above 81% in one year. It also demonstrates software-comparable performance in wireless communication symbol detection tasks. In the two tasks tested, Xpikeformer achieved 17.817.817.817.819.2×19.2$\times$19.2 × energy reduction compared to the state-of-the-art digital ANN transformer accelerator, a 5.05.05.05.07.1×7.1$\times$7.1 × energy reduction compared to the hybrid implementation of ANN, and a 5.95.95.95.96.8×6.8$\times$6.8 × energy reduction compared to the fully digital implementation of spiking transformers on general-purpose devices. Xpikeformer also achieves a 1.53×1.53$\times$1.53 × speedup than traditional ANN on GPU and a 12.0×12.0$\times$12.0 × speedup compared to the GPU implementation of spiking transformers.
In this work, we adopted Neurosim’s crossbar assumption to facilitate energy and latency evaluation. Nevertheless, this approach may not yield the best performance in terms of energy efficiency. Practical implementation of memristor crossbars usually adopts a different input strategy by using the input spikes to drive the conductivity of devices through word lines and using inverted voltage spikes and differential-input ADCs to represent positive and negative values in the crossbar [54].
Future work may include exploring more energy-efficient crossbar designs that better align with state-of-the-art implementations. This could involve investigating alternative input strategies and circuit designs that optimize the use of inverted voltage spikes and differential-input ADCs. Additionally, the integration of advanced memristor technologies and the development of new materials for memristors, e.g., 2-D material, could further enhance the energy efficiency and performance of neuromorphic computing systems [55, 56]. Evaluating the impact of these practical crossbar implementations on the overall system performance and conducting comprehensive real-world testing will be crucial steps in advancing the applicability of Xpikeformer and similar models in energy-constrained environments.

### v2

@@ VII-A3 Results
Fig. 9 presents the breakdown of Xpikeformer computational energy consumption. The majority of the computing energy is consumed by the AIMC engine, accounting for 78.4% of the total computational energy. The SSA engine contributes 18.9%, while other components, such as residual units, collectively account for 2.7%. A more detailed breakdown of the AIMC engine’s energy consumption is shown on the right side of Fig. 9. Notably, the ADC consumes only 2.0% of the AIMC energy. This efficiency is achieved by sharing ADCs across multiple read-out channels through multiplexing (Xpikeformer uses a sharing ratio of 8, balancing energy efficiency and latency), which significantly reduces static power consumption that would otherwise be incurred by using dedicated ADCs for each channel. We also observe that periphery circuits (decoders, multiplexers, switch matrices, buffers, etc.) consume 85.9% of the AIMC energy. This is because ADC multiplexing requires a dedicated decoder and control logic, making these components relatively more energy-intensive than crossbar itself. Additionally, the SA mapping strategy introduces additional control overhead between shared SRAM and local SA buffers. This overhead is unavoidable in this work due to the crossbar size limitations discussed in Section IV-A. Finally, accumulation processes, including differential adders and LIF units, which involve digital operations, contribute 12.1% of AIMC energy.
VII-B Latency and Area Evaluation
The latency breakdown of Xpikeformer is shown in Fig. 10. More than 92% of the latency is attributed to peripheral circuit operations, as these handle global data movement, routing, and control tasks, which are inherently sequential and cannot be parallelized like core computations. Additionally, decoding, buffering, and interconnect operations scale with model size, further contributing to the overhead. In contrast, the actual AIMC computing latency is negligible, accounting for only 0.3% of the total latency. The SSA computation contributes only 2.0% of the overall latency.
We compare the per-inference latency of Xpikeformer with two same-sized baseline implementations: an ANN-based transformer and spiking transformer as in [15], both deployed on a GPU (Nvidia RTX A2000), adopting the minimum spiking encoding lengths as in Section VI-A. Results are shown in Fig. 10. Despite the extra temporal dimension, Xpikeformer achieves a 2.18×2.18$\times$2.18 × speedup over the GPU implementation of ANN-based transformers, benefiting from fast in-memory MVM and the parallel, streamlined SSA design. Meanwhile, the spiking transformer suffers from poor GPU efficiency due to precision mismatches and lack of hardware optimization. In contrast, Xpikeformer delivers a 6.85$\times$ speedup over it.
Figure 10: (a) Latency breakdown of Xpikeformer and (b) Latency comparison with ANN and SNN transformers on GPU.
The area of Xpikeformer’s computing engines is estimated using the DNN+NeuroSim V1.4 framework and the Cadence synthesis tool at the 45 nm technology node. The total chip area of Xpikeformer is 784 mm2, with the majority (76.5%) occupied by periphery circuits and interconnections. The AIMC computing core, which includes crossbars, ADCs, and accumulation units, accounts for 11.5% of the area. The SSA engine, being more area-efficient than traditional ALUs due to its use of lightweight logic gates as computing units, contributes 12% of the total area.
VII-C Comparison With SOTA Hardware Accelerators
To the best of the authors’ knowledge, Xpikeformer is the first hardware accelerator designed for spiking transformers. We compare Xpikeformer with two SOTA ASIC accelerators for ANN-based transformers, summarized in Table VI.
SwiftTron [34] is a fully digital ASIC accelerator designed for quantized transformers. It replaces traditional floating-point ALUs with fixed-point computing units and approximates nonlinear operations using second-order polynomials. The reported chip area of SwiftTron is 2.9$\times$ smaller than that of Xpikeformer. However, this reduction is achieved by reusing computational resources through time multiplexing, which introduces additional energy consumption and latency due to frequent parameter reads during inference. Benchmarked on the same task, Xpikeformer achieves similar latency while consuming 13$\times$ less energy by leveraging AIMC for feed-forward layers and lightweight logic-gate-based computations for the attention mechanism.
X-Former [24] is an in-memory accelerator for ANN transformers that employs AIMC crossbars for feed-forward layers and DIMC for attention. Note that X-Former architecture uses 1-bit ReRAM for weight storage, which translates to lower weight storage density compared to Xpikeformer’s multi-bit PCM devices, potentially increasing the overall area for AIMC implementation. For MHSA computation, X-Former employs DIMC units based on SRAM arrays, which require writing $Q$ and $V$ matrices into SRAM crossbars during inference, and storing intermediate results in the memory, leading to additional latency. In contrast, Xpikeformer leverages the SSA engine to eliminate online memory-write operations and intermediate storage while also avoiding the complex peripherals associated with DIMC. Overall, Xpikeformer achieves a 6$\times$ reduction in inference energy and a 1.9$\times$ speedup over X-Former.
TABLE VI: Comparison With SOTA Accelerators
Performance Metric [34] [24] Xpikeformer
Paradigm ANN ANN SNN
MAC Implementation Digital ALU ReRAM-AIMC PCM-AIMC
MHSA Implementation Digital ALU DIMC SSA
Technology 65 nm 32 nm 45 nm
Weight INT8 INT8 (Equiv.) INT5 (Equiv.)
Activation Precision INT8/32 INT8 Multi-Step Binary
Frequency 143 MHz 200 MHz 200 MHz
Area (mm2) 273.0 784
Energy/Inference* (mJ) 3.97 2.04 0.30
Latency/Inference* (ms) 2.26 4.13 2.18
  • * Energy and latency results have been normalized to a consistent benchmark task (ImageNet-1K classification on ViT-8-768 with a patch size of $16\times 16$) to ensure fair comparison across different architectures. For [34], we assume their original chip area is fixed and scale the overall latency proportionally to the increase in task size. For [24], we assume that AIMC resources are sufficient to accommodate all model parameters, while the DIMC resources remain fixed as originally reported, with attention latency scaled up accordingly.
  • The latency results for [24] and Xpikeformer account for both computational latency and data movement from memory to computing cores.
VII-D Scalability of Xpikeformer
For the SSA engine, separating SSA tiles from the LFSR array improves area efficiency by centralizing PRN generation in the shared LFSR array, leaving only basic logic gates and bit-wise counters in the tiles. In contrast, traditional fixed-point ALU arrays with adders and multipliers incur higher area overhead. The SSA engine’s 1-bit I/O buses, with no intermediate storage or operations, ensure high data-dimensional flexibility. Its stateless, modular design supports both head-wise and layer-wise reuse, minimizing resource duplication and enabling efficient scaling to larger attention heads and additional layers without significant hardware overhead.
Like most NVM-based architectures, the AIMC engine in Xpikeformer scales spatially to accommodate larger weight matrices. By leveraging the multi-level programming capability of PCM devices, it achieves a 3$\times$ reduction in required memory devices compared to 1-bit ReRAM cells [24]. With similar planar cell areas ($4F^{2}$$8F^{2}$) to ReRAM, PCM crossbars provide denser weight storage, improving efficiency in device count and layout area for large-scale models. The row-block-wise mapping strategy further enhances memory utilization and weight distribution. For very-large models, Xpikeformer can scale further by connecting multiple nodes via chip-to-chip interconnects.
@@ VIII Conclusion
In this paper, we introduced Xpikeformer, a hybrid analog-digital hardware architecture designed to accelerate spiking neural network (SNN)-based transformer models. Xpikeformer integrates analog in-memory computing (AIMC) for feedforward and fully connected layers and a stochastic spiking attention (SSA) engine for efficient attention computation, mitigating the inefficiencies of deploying spiking transformers on general-purpose platforms.
Trained from scratch, spiking transformer models running on Xpikeformer achieve competitive accuracy on ImageNet-1K classification and wireless communication symbol detection tasks while maintaining high accuracy for over a year. Although the non-reusable nature of AIMC introduces an area tradeoff compared to non-AIMC digital accelerators implementing time multiplexing, Xpikeformer achieves 13$\times$ reduction in energy consumption at approximately the same throughput as the SOTA digital accelerator for ANN-based transformers.
In this work, we adopted NeuroSim’s crossbar assumptions to facilitate energy, latency, and area evaluations. However, this approach may not fully capture the performance of SOTA AIMC technology. Practical implementations of NVM crossbars often adopt different input strategies, such as using input spikes to drive device conductivity through word lines and employing inverted voltage spikes and differential-input ADCs to represent positive and negative values in the crossbar [57]. Future work will explore more energy-efficient NVM crossbar designs, advanced materials like 2D semiconductors, and real-world testing to further enhance Xpikeformer’s efficiency and applicability in energy-constrained environments.